Particle A of charge 3.50 x 10-4 C is at the origin, particle 8 of charge -6.0s x 10-4 C is at (4.26 m, O) and particle C of charge 1.1 a) What is the x-component of the electric force exerted by A on C (b) What is the y-component of the force exerted by A on c? (c) Find the magnitude of the force exerted by B on C (d) Calculate the x-component of the force exerted by B on C. (e) Calculate the y component of the force exerted by B on C D Sum the two x-components to obtain the resultant x-component of the electric force acting on c (g) Repeat part (f) for the y-component. (h) Find the magnitude and direction of the resultant electric force acting on C. magnitude counterclockwise from the +x-axis
Answer #1

FACy FBC

qA = 3.5*10^-4 C

qB = -6.05*10^-4 C

qC = 1.15*10^-4 C


AB = 4.26 m

AC = 3.56 m

BC = sqrt(AC^2+BC^2) = sqrt(3.56^2 + 4.26^2) = 5.55 m

part (a)

EACx = 0


part (b)


FACy = k*qA*qC/AC^2 =


FACy = 9*10^9*3.5*10^-4*1.15*10^-4/3.56^2


FACy = 28.6 N


part( c)

FBC = k*qB*qC/BC^2

FBc = 9*10^9*6.05*10^-4*1.15*10^-4/5.55^2

FBC = 20.3 N

part D

FBCx = FB*sinC = FBC*AB/BC = 20.3*(4.26/5.55) = 15.6 N


part (E)


FBy = -FB*cosC = -FBC*AC/BC = -20.3*(3.56/5.55) = -13.02 N


part (F)

Fcx = FACx + FBCx = 0 + 15.6 = 15.6 N


(g)

FCy = FACy + FBCy = 28.6 - 13.02 = 15.58 N


(h)

magnitude = sqrt(15.6^2+15.58^2) = 22.05 N


direction = tan^-1(Fcy/Fcx)


direction = 44.9